# ON THE AVERAGE ULTRAVIOLET EMISSION LINE SPECTRA OF HIGH-REDSHIFT GALAXIES: HOT AND COLD, CARBON-POOR, NITROGEN-MODEST, AND OOZING IONIZING PHOTONS

### v1

@@ 6.3.1 Carbon
Four of the stacks shows log(C/O) , with uncertainties as low as dex for the highest (youngest) starbursts, where the S/N ratio of the UV lines is highest. This uncertainty increases for the lower stacks. The central bin in scatters upwards to log(C/O) , but is not significantly different when accounting for the uncertainty. These measurement are broadly consistent with those in the literature for low metallicity galaxies at low-redshift, and also some observed at using recent JWST spectroscopy (Figure 8). Our C/O abundances also agree with the average value by Hu et al. (2024) from the stacked analysis of 63 JWST galaxies at , using higher resolution data.
The C/O ratio should increase with time during the evolution of a galaxy: it starts low (log(C/O) ), where the ratio simply reflects the yields of core-collapse supernova (CCSN). CCSN explode almost immediately after an episode of star formation, and produce the overwhelming majority of oxygen in our universe, with yields that are reasonable certain. However, the bulk of carbon is produced by low and intermediate mass stars, as they assent the Asymptotic Giant Branch (AGB) on timescales of 100 Myr to several Gyr. Thus at time goes on, the (C/O) ratio increases from some plateau value as proportionally more carbon than oxygen is delivered to the ISM.
Our measurements of the oxygen abundance (Figure 8) absolutely show the expected increase in , which grows from at age of 3 Myr to at age of 60 Myr (See Figure 6 for ages). Interesting, however, our (C/O) measurements do not capture the expected buildup of carbon: log(C/O) appears fixed at for all metallicities where it can be measured. A large number of literature points from both low- and high-redshifts are over-plotted to illustrate the known variations in C/O from other samples. This is likely because the light-weighted age we recover (e.g. Figure 6 extend only to ages of Myr for stacks in which C/O can be measured – these ages are far from sufficient to capture the major fraction of carbon production, which is highest some 500 Myr after major episodes of star formation (Mattsson, 2010). Of course star formation history adds a more nuanced view than the simple quantity of ‘age’, and we have made no efforts to reconstruct more detailed SFHs in this paper, nevertheless, if a significant fraction of the stellar population has formed in the last half billion years, which seems plausible for strong line-emitting galaxies at on average, then then it is hard to avoid the under-production of carbon with respect to oxygen.
@@ 6.3.2 Nitrogen
We show the N/O ratios in Figure 9. Recent papers by Topping et al. (2024) and Schaerer et al. (2024) show compilations of N/O measurements vs. 12+log(O/H) for a sample of local galaxies including CLASSY and the starburst observed by (Senchyna et al., 2022), as well as a handful of individually detected ‘nitrogen-loud’ galaxies at high-. We show these systems as literature points in Figure 9, together with the measurements in our stacks. In this version of the figure, we provide measurements using basic ionization correction factors, for the sake of comparison with the literature points. Our stacks cover the same range in oxygen abundance as the individual sources. Our estimates of log(N/O) fall in the range to 0.0 in the stacks for which N iii] and/or N iv] can be detected. This is in excellent agreement with the measurements made at high-, as well as local starburst Mrk 996 which have broadly similar oxygen abundances. Within the uncertainties, we do not find evolution of N/O with 12+log(O/H). Since our estimates use similar methods to those of the galaxies, we must conclude that if these individually detected objects show ‘exotic conditions’, then the same exotic conditions must be found in the average galaxy over the redshift range of this survey.
Regarding the cause of the strong nitrogen enrichment in galaxies, we can now reject explanations in order of increasing implausibility. This means we can safely rule out the possibility of the enhanced nitrogen emission being associated with tidal disruption events, since it is implausible that the median galaxy hosts one ongoing TDE. For similar reasons, it would be a bold claim that AGN are the culprits: while they have been recently discovered in very high numbers at these high redshifts (Maiolino et al., 2023; Harikane et al., 2023b; Hayes et al., 2024; Kocevski et al., 2024; Matthee et al., 2024) there is an enormous difference between an AGN occupation fraction of a few vs. 100 percent. Moreover, stronger emission line diagnostics in the UV and optical do not show substantial evidence for AGN in these stacked spectra.
@@ 6.4 Production and Emission of Ionizing Radiation
As shown in Figure 10, increases by a factor of across our stacks, from log(/[Hz erg to 25.65. These estimates are perfectly in line with those reported by Stefanon et al. (2022) and Simmonds et al. (2024a) for emission line galaxies at similar redshifts (and indeed include some of the same galaxies used by Simmonds et al. 2024a). They slightly exceed estimates from Simmonds et al. (2024b) for the broader sample of JADES galaxies, and likely for reasons outlined therein: our demand for measured emission lines (for accurate spectroscopic redshifts) biases the sample towards higher values of . The higher end of our range is consistent with the value of 25.8 reported by Atek et al. (2024) for strongly gravitationally lensed galaxies: our least luminous stack (highest ) has , which is still a factor of more luminous than those in the Atek et al sample and consistent with the expectation that increases further towards lower masses and luminosities (although see also Simmonds et al. 2024b). Even if we cannot provide evidence for the star formation in the high- galaxies being bursty, the stellar populations of these galaxies are consistent with ages of just a few Myr (Figure 6) and these values of are sufficient to reionize the Universe if the escape fractions are high enough.
The remaining three panels of Figure 10 all shows that the escape fractions of ionizing radiation are probably high. Firstly, even the plot of (left) shows the high- sources produce more ionizing photons per UV luminosity than their low- counterparts: we compare the JWST measurements with those from the Low Redshift Lyman Continuum Survey (LzLCS Flury et al., 2022b, a). Recent modeling results in LzLCS show that the highest values are reported for the youngest galaxies, implicating the radiative feedback that occurs only at very young ages (Carr et al., 2024; Flury et al., 2024). Radiative feedback in these high- galaxies is likely to be higher still, further elevating LyC emission.
@@
We also show the [S ii]/H vs. [O iii]/H BPT diagram in the lower left panel of the same Figure. This diagnostic is motivated by Alexandroff et al. (2015) and Wang et al. (2019, see also ): the ionization potential range of singly ionized sulphur is 10.4–23.3 eV causes the [S ii] emitting regions to lie towards the outer edge of the Strömgren sphere. Truncated H ii regions, therefore, will show a deficit of [S ii] emission compared to line ratios that trace more highly ionized gas (e.g. the [O iii]/H ratio on the x-axis). The locus of SDSS galaxies is illustrated in this plot, together with the best fit to these objects from Wang et al. (2021). The first three of our high-z stacks lie close to the SDSS ridge line and Wang et al. (2021) prescription, albeit at the very high end in the [O iii]/H plane. They occupy a similar range in the parameter space to the more extreme half of galaxies in LzLCS in [O iii]/H space. However two the highest stacks fall far from the Wang et al. line, by 0.36 and 0.89 dex, which flow through the most extreme deficits seen in the LzLCS program. The highest [S ii] deficit is larger than any of the objects in LzLCS ([S ii] = 0.66). As with the Mg ii discussion above, correlations between [S ii] and exhibit too much spread to use for quantitative estimates, but we note that if we take the LzLCS objects for which [S ii] exceeds 0.3 dex then the mean reaches a significant value of 12%.
We finally show the variation of the UV slope () and O32 ratio in the lower right panel of Figure 10. There is a monotonic bluening of the continuum color and increase in the ionization parameter across the sequence of . Again, these tell the story of stellar populations becoming younger (verified by the stellar population fits in Figure 6), less dust obscured (verified by the H/H ratios), and lower metallicity (not verified by stellar metallicities, but demonstrated in the gas phase by ; Figure 8). The values measured in the highest stacks fall into the range where Chisholm et al. (2022) estimate to be in the 5–10% range. These two stacked spectra show O32 ratios between 10 and 15, at which range Flury et al. (2022a) would have found a LyC detection rate above 50% for low- analog galaxies. Combined with the high ionizing photon production efficiencies discussed above, we expect these samples to be very bright in the Lyman continuum.
@@ 7 Summary and Conclusions
We have studied ostensibly star-forming galaxies at redshifts between 4 and 10 observed with JWST/NIRSpec in the low-resolution PRISM mode. We executed a stacking analysis, dividing the galaxies into five sub-samples, sorted by their equivalent width in the [O iii]5008 line, reaching total integration times of million seconds per spectral pixel. This provides the depth needed to reveal faint ultraviolet emission lines in [C ii], C iii], C iv, He ii, O iii], N iii], N iv] and Mg ii which are not regularly observed in JWST spectra of high- galaxies. We also detect, at very high signal-to-noise, a complete suite of optical emission lines at wavelengths between [O ii] and [S ii].
@@
We measure the luminosity and restframe equivalent width of these emission lines, paying particular attention to the UV. We show that all features in the UV increase in strength as a function of [O iii] strength, and much of this evolution can naturally be attributed to the decreasing average age of the stellar population with increasing . We place the measurements from our 5 individual stacks on UV diagnostic diagrams that show line ratios and EWs, concluding they are predominantly star-forming, although we do find evidence that the spectral slope of the extreme ultraviolet becomes steeper towards younger galaxies.
@@
We measure the interstellar gas temperatures using the [O iii]1666/5008 line ratios, deriving temperatures in the range 17,000 to 22,000 Kelvin. However our very deep stacked spectra also well-expose the Balmer break in the optical, which is clearly seen and allows an independent measurement of the average electron temperature. The is % lower than that derived from the [O iii] ratio, a difference that is larger than would be expected from typical temperature differences from the O++ and O+ zones.
@@
We use these temperatures to measure accurate abundances for oxygen, carbon, and nitrogen, finding oxygen abundances in the range . Consistent with other studies at the same redshift, we find carbon/oxygen ratios at log(C/O); however less consistently, we find no evidence of an increase in C/O with O/H, despite the more metallic galaxies being manifestly more evolved. This may plausibly be explained by us catching galaxies in a strong starburst phase (demanding [O iii] detections for redshifts), but still question why previous studies would not find the same lack of evolution.
@@
Nitrogen appears to be somewhat enriched, and we find values of log(N/O) in the three highest for with N iii] and N iv] are detected. This is consistent with recent estimates in individual galaxies. The individual galaxies have been discussed extensively in many recent papers, and attributed to hidden AGN, tidal disruption event, or the very massive stars that may signpost the formation of proto-globular clusters. Regardless of the exact origin of these emission lines, we believe the same processes are running at a lower level in the average galaxy at these redshifts.
@@
We investigate various signatures of the production and escape of Lyman continuum emission, showing that the ionizing photon production efficiency varies between log(/[Hz erg-1]) and 25.65. Concerning escape fractions we focus uoon the Mg ii and [S ii] emission lines, [O iii]/[O ii] ratios and UV continuum color. We compute the Mg ii escape fraction, and show it varies between zero (and , as Mg ii is seen in absorption in the weakest emission stacks) and almost 1, with for the two strongest stacks. In the same two stacks the [S ii]/H ratio falls rapidly away from the locus of local galaxies, O32 ratios increase to above 10, and UV slopes become bluer than . All these diagnostics have been connected with high in the local Universe, and we expect similar phenomena are at play in these reionization era galaxies.

### v2

@@ 6.3.1 Carbon
Four of the stacked spectra show log(C/O) , with uncertainties as low as dex for the highest (youngest) starbursts, where the S/N ratio of the UV lines is highest. This uncertainty increases for the lower values of where the lines become weaker. The central bin in scatters upwards to log(C/O) , but is not significantly different when accounting for the uncertainty. These measurement are broadly consistent with those in the literature for low metallicity galaxies at low-redshift, and also some observed at using recent JWST spectroscopy (Figure 8). Our C/O abundances also agree with the average value by Hu et al. (2024) from the stacked analysis of 63 JWST galaxies at , using higher resolution data.
The C/O ratio should increase with time during the evolution of a galaxy: it starts low (log(C/O) ), where the ratio simply reflects the yields of core-collapse supernova (CCSN). CCSN explode almost immediately after an episode of star formation, and produce the overwhelming majority of oxygen in our universe, with yields that are reasonably certain. However, the bulk of carbon is produced by low and intermediate mass stars, as they ascend the Asymptotic Giant Branch (AGB) on timescales of 100 Myr to several Gyr. Thus at time goes on, the (C/O) ratio increases from some plateau value as proportionally more carbon than oxygen is delivered to the ISM.
Our measurements of the oxygen abundance (Figure 8) absolutely show the expected increase in , which grows from at age of 3 Myr to at age of 60 Myr (See Figure 6 for ages). Interesting, however, our (C/O) measurements do not capture the expected buildup of carbon: log(C/O) appears fixed at for all metallicities (ages) where it can be measured. A large number of literature points from both low- and high-redshifts are over-plotted to illustrate the known variations in C/O from other samples. This is likely because the light-weighted age we recover (e.g. Figure 6 extend only to ages of Myr for stacks in which C/O can be measured – these ages are far from sufficient to capture the major fraction of carbon production, which is highest some 500 Myr after major episodes of star formation (Mattsson, 2010). Of course star formation history adds a more nuanced view than the simple quantity of ‘age’, and we have made no efforts to reconstruct more detailed SFHs in this paper, nevertheless, if a significant fraction of the stellar population has formed in the last half billion years, which seems plausible for strong line-emitting galaxies at on average, then then it is hard to avoid the under-production of carbon with respect to oxygen.
@@ 6.3.2 Nitrogen
We show the N/O ratios in Figure 9. Recent papers by Topping et al. (2024) and Schaerer et al. (2024) show compilations of N/O measurements vs. 12+log(O/H) for a sample of local galaxies including CLASSY and the starburst observed by (Senchyna et al., 2022), as well as a handful of individually detected ‘nitrogen-loud’ galaxies at high-. We show these systems as literature points in Figure 9, together with the measurements in our stacks. In the left panel of the figure, we provide measurements using basic ionization correction factors, for the sake of comparison with the literature points. Our stacks cover the same range in oxygen abundance as the individual sources. Our estimates of log(N/O) fall in the range to 0.0 in the stacks for which N iii] and/or N iv] can be detected. This is in excellent agreement with the measurements made at high-, as well as local starburst Mrk 996 which have broadly similar oxygen abundances. Within the uncertainties, we do not find evolution of N/O with 12+log(O/H) although the estimate values do correlate. Since our estimates use similar methods to those of the galaxies, we must conclude that if these individually detected objects show ‘exotic conditions’, then the same exotic conditions must be found in the average galaxy over the redshift range of this survey.
Regarding the cause of the strong nitrogen enrichment in galaxies, we can now reject explanations in order of increasing implausibility. This means we can safely rule out the possibility of the enhanced nitrogen emission being associated with tidal disruption events, since it is implausible that the median galaxy hosts one ongoing TDE. For similar reasons, it would be a bold claim that AGN are the culprits: while they have been recently discovered in very high numbers at these high redshifts (Maiolino et al., 2023; Harikane et al., 2023b; Hayes et al., 2024; Kocevski et al., 2024; Matthee et al., 2024) there is an enormous difference between an AGN occupation fraction of a few percent and the % needed to be visible in a median stack. Moreover, stronger emission line diagnostics in the UV and optical do not show substantial evidence for AGN in these stacked spectra, although we note again that the [C ii]2326 line is only visible in the stack with the highest N/O.
@@ 6.4 Production and Emission of Ionizing Radiation
As shown in Figure 10, increases by a factor of across our stacks, from log(/[Hz erg to 25.65. These estimates are perfectly in line with those reported by Stefanon et al. (2022) and Simmonds et al. (2024a) for emission line galaxies at similar redshifts (and indeed include some of the same galaxies used by Simmonds et al. 2024a). Our estimates of slightly exceed those from Simmonds et al. (2024b) for the broader sample of JADES galaxies, and likely for reasons outlined therein: our demand for measured emission lines (for accurate spectroscopic redshifts) biases the sample towards higher values of . The higher end of our range is consistent with the value of 25.8 reported by Atek et al. (2024) for strongly gravitationally lensed galaxies: our least luminous stack (highest ) has , which is still a factor of more luminous than those in the Atek et al sample and consistent with the expectation that increases further towards lower masses and luminosities (although see also Simmonds et al. 2024b). Even if we cannot provide evidence for the star formation in the high- galaxies being bursty, the stellar populations of these galaxies are consistent with ages of just a few Myr (Figure 6) and these values of are sufficient to reionize the Universe if the escape fractions are high enough.
The remaining three panels of Figure 10 all show that the escape fractions of ionizing radiation are probably high. Firstly, even the plot of (left) shows the high- sources produce more ionizing photons per UV luminosity than their low- counterparts: we compare the JWST measurements with those from the Low Redshift Lyman Continuum Survey (LzLCS Flury et al., 2022b, a). Recent modeling results in LzLCS show that the highest values are reported for the youngest galaxies, implicating the radiative feedback that occurs only at very young ages (Carr et al., 2024; Flury et al., 2024). Radiative feedback in these high- galaxies is likely to be higher still, further elevating LyC emission.
@@
We also show the [S ii]/H vs. [O iii]/H BPT diagram in the lower left panel of the same Figure. This diagnostic is motivated by Alexandroff et al. (2015) and Wang et al. (2019, see also ): the ionization potential range of singly ionized sulfur is 10.4–23.3 eV, causing the [S ii] emitting regions to lie towards the outer edge of the Strömgren sphere. Truncated H ii regions, therefore, will show a deficit of [S ii] emission compared to line ratios that trace more highly ionized gas (e.g. the [O iii]/H ratio on the x-axis). The locus of SDSS galaxies is illustrated in this plot, together with the best fit to these objects from Wang et al. (2021). The first three of our high-z stacks lie close to the SDSS ridge line and Wang et al. (2021) prescription, albeit at the very high end in the [O iii]/H plane. They occupy a similar range in the parameter space to the more extreme half of galaxies in LzLCS in [O iii]/H space. However two the highest stacks fall far from the Wang et al. line, by 0.36 and 0.89 dex, where the average galaxy exceeds the most extreme deficit seen in the LzLCS program ([S ii] = 0.66). As with the Mg ii discussion above, correlations between [S ii] and exhibit too much spread to use for quantitative estimates, but we note that if we take the LzLCS objects for which [S ii] exceeds 0.3 dex then the mean reaches a significant value of 12%.
We finally show the variation of the UV slope () and O32 ratio in the lower right panel of Figure 10. There is a monotonic bluening of the continuum color and increase in the ionization parameter across the sequence of . Again, these tell the story of stellar populations becoming younger (verified by the spectral fits in Figure 6), less dust obscured (verified by the H/H ratios), and lower metallicity (not verified by stellar metallicities, but demonstrated in the gas phase by ; Figure 8). The values measured in the highest stacks fall into the range where Chisholm et al. (2022) estimate to be in the 5–10% range. These two stacked spectra show O32 ratios between 10 and 15, at which range Flury et al. (2022a) would have found a LyC detection rate above 50% for low- analog galaxies. Combined with the high ionizing photon production efficiencies discussed above, we expect these samples to be very bright in the Lyman continuum.
@@ 7 Summary and Conclusions
We have studied ostensibly star-forming galaxies at redshifts between 4 and 10 that were observed with JWST/NIRSpec in the low-resolution PRISM mode. We performed stacking analyses on the spectra, dividing galaxies into five sub-samples sorted by their equivalent width in the [O iii]5008 line. These combined spectra reach total integration times of million seconds per spectral pixel. This provides the depth needed to reveal faint ultraviolet emission lines in [C ii], C iii], C iv, He ii, O iii], N iii], N iv] and Mg ii which are not regularly observed in JWST spectra of high- galaxies. We also detect, at very high signal-to-noise, a complete suite of optical emission lines at wavelengths between [O ii] and [S ii].
@@
We measure the luminosity and restframe equivalent width of these emission lines, paying particular attention to the UV. We show that all features in the UV increase in strength as a function of [O iii] EW, and much of this evolution can naturally be attributed to the decreasing average age of the stellar population with increasing . We place the measurements from our five individual stacks on UV diagnostic diagrams that show line ratios and EWs, concluding they are predominantly star-forming, although we do find evidence that the spectral slope of the extreme ultraviolet becomes steeper towards younger galaxies.
@@
We measure the interstellar gas temperatures using the [O iii]5008/1666 line ratios, deriving temperatures in the range 17,000 to 22,000 Kelvin. However our very deep stacked spectra also well-expose the nebular Balmer jump in the optical, which is clearly seen and allows an independent measurement of the average electron temperature. The is % lower than that derived from the [O iii] ratio, a difference that is larger than would be expected from typical temperature differences from the O++ and O+ zones.
@@
We use these temperatures to measure accurate abundances for oxygen, carbon, and nitrogen, finding oxygen abundances in the range . Consistent with other studies at the same redshift, we find carbon/oxygen ratios at log(C/O); however less consistently, we find no evidence of an increase in C/O with O/H, despite the stellar populations of the more metallic galaxies also being shown to be more evolved. This may plausibly be explained by us catching galaxies in a strong starburst phase (demanding [O iii] detections for redshifts), but does not erase the question of why previous studies do not find the same lack of evolution.
@@
Nitrogen appears to be somewhat enriched, and we find values of log(N/O) in the three highest stacks, where N iii] and N iv] are detected. This is consistent with recent estimates in individual galaxies. The individual galaxies have been discussed extensively in many recent papers, and attributed to hidden AGN, tidal disruption event, or the very massive stars that may signpost the formation of proto-globular clusters. Regardless of the exact origin of these emission lines, we believe the same processes are running at a lower level in the average galaxy at these redshifts. We expect this is due to stellar activity, but cannot entirely rule out that it is relate to the presence of SMBHs.
@@
We investigate various signatures of the production and escape of Lyman continuum emission, showing that the ionizing photon production efficiency varies between log(/[Hz erg-1]) and 25.65. Concerning escape fractions we focus upon the Mg ii and [S ii] emission lines, [O iii]/[O ii] ratios and UV continuum color. We compute the Mg ii escape fraction, and show it varies between zero (and , as Mg ii is seen in absorption in the weakest emission stacks) and almost 1, with for the two strongest stacks. In the same two stacks the [S ii]/H ratio falls rapidly away from the locus of local galaxies, O32 ratios increase to above 10, and UV slopes become bluer than . All these diagnostics have been connected with high in the local Universe, and we confidently believe the highest galaxies in our sample are strong LyC-leaking galaxies.